OpenStudy (anonymous):

find the sum . 6/18 + 3/19 = 9/37 ?

OpenStudy (anonymous):

{(6*19)/(18*19)}+{(3*18)/(19*18)=(1/3)+(3/19) (19/57)+(9/57)=28/57